OpenStudy (anonymous):

The spectrophotometer really measures the percent of light that is transmitted through the solution. The instrument then converts %T (transmittance) into absorbance by using the equation you determined in the prelab section. If the absorbance is 0.85, what is the percent light transmitted through the color sample at this collected wavelength?

OpenStudy (anonymous):

the equation is y=0.073487(conc)+0.024531

OpenStudy (aaronq):

what is the y variable in the equation? you can use: \(Abs = -log (\dfrac{\%T}{100})\) you should double check that formula in your notes, i don't know if it's 100% correct

OpenStudy (anonymous):

my bad wrong one that is for something else

OpenStudy (aaronq):

so the equation is actually: \(Abs=-log(T)\) you just need to convert that to a percent after

OpenStudy (anonymous):

\[-\log_{.85/100} \]

OpenStudy (anonymous):

-log(85%)

OpenStudy (anonymous):

is the answer 1.92941

OpenStudy (aaronq):

i don't think that's right.

OpenStudy (aaronq):

it shouldn't be over 100%

OpenStudy (anonymous):

i g-0.07058 after redoing and changing it to a decimal

OpenStudy (anonymous):

-log(1-0.85)

OpenStudy (aaronq):

absorbance is on the other side of the equal sign

OpenStudy (anonymous):

-log(0.85)

OpenStudy (aaronq):

no absorbance is on the other side, you're plugging into where transmittance goes

OpenStudy (anonymous):

I do not understand

OpenStudy (aaronq):

\(\huge \color{red}{Abs}=-log(T)\)

OpenStudy (anonymous):

can you show me what you mean

OpenStudy (aaronq):

0.85=-log(T)

OpenStudy (anonymous):

okay a

OpenStudy (anonymous):

so I just need to figure out how to find T right

OpenStudy (aaronq):

then convert it to a percent

OpenStudy (anonymous):

so do I multiple 0.85 by 100 to get t then

OpenStudy (aaronq):

no, once you find T, multiply by 100%